A Body Makes three equal displacements each with velocities V1,V2 And V3.What Is The Average Velocity  Of The Body

Grade:11

1 Answers

Rishabh Gaikwad
6 years ago
Since displacements are equal
Therefore, S1 = S2 = S3 = 1
Velocity = displacement/time
V1 = S1/T1 => T1 = S1/V1 => 1/V1
V2 = S2/T2 => T2 = S2/V2 => 1/V2
V3 = S3/T3 => T3 = S3/V3 => 1/V3
 
Average velocity = Total Displacement/Total Time
                         = S1 + S2 + S3/T1 + T2 + T3
                         = 1 + 1 + 1/1/V1 + 1/V2 + 1/V3
                         = 3/1/V1 +1/V2 +1/V3
I hope this helps you.